# Pushing the Accuracy Limit of Foundation Neural Network Models with Quantum Monte Carlo Forces and Path Integrals

### v2

@@ Acknowledgment
QMC, sCI calculations, DFT computations on Aurora and some molecular dynamics simulation performed in this research used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (AB, KG, JPP). DFT Calculations on Polaris used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (RCC, JTK and YL).This research has used resources (Aurora and Polaris machines) of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. RCC (QMC forces algorithms), JTK (QMC workflow engine) and YL (QMCPACK code development) were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. CK and VA were supported by the Computational Chemical Sciences Program, under Award DE-SC0023382, funded by the U.S. Department of Energy, Office of Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division. This work was also made possible thanks to funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 810367), project EMC2 (JPP). We acknowledge EuroHPC Joint Undertaking for awarding the project IDs EHPC-DEV-2024D07-044 (E. P.) and EHPC-AI-2024A04-085 (E. P.) access to Leonardo at CINECA, Italy for DFT computations. The learning of the FeNNix-Bio1 model and some molecular dynamics simulations were performed at IDRIS, GENCI (Jean Zay machine, France) on grants No. A0150712052 (J.-P.P.) and grant GC010815453 (Grand Challenge H100 Jean Zay, J.-P.P.).

### v3

@@ Acknowledgment
QMC, sCI calculations, DFT computations on Aurora and some MD simulations performed in this research used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (AB, KG, JPP). DFT Calculations on Polaris used an award of computer time provided by the Innovative and Novel Computational Impact on Theory and Experiment (INCITE) program (RCC, JTK and YL).This research has used resources (Aurora and Polaris machines) of the Argonne Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC02-06CH11357. RCC (QMC forces algorithms), JTK (QMC workflow engine) and YL (QMCPACK code development) were supported by the U.S. Department of Energy, Office of Science, Basic Energy Sciences, Materials Sciences and Engineering Division, as part of the Computational Materials Sciences Program and Center for Predictive Simulation of Functional Materials. CK and VA were supported by the Computational Chemical Sciences Program, under Award DE-SC0023382, funded by the U.S. Department of Energy, Office of Basic Energy Sciences, Chemical Sciences, Geosciences, and Biosciences Division. This work was also made possible thanks to funding from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No 810367), project EMC2 (JPP). We acknowledge EuroHPC Joint Undertaking for awarding the project IDs EHPC-DEV-2024D07-044 (E. P.) and EHPC-AI-2024A04-085 (E. P.) access to Leonardo at CINECA, Italy for DFT computations. The learning of the FeNNix-Bio1 model and some MD simulations were performed at IDRIS, GENCI (Jean Zay machine, France) on grants No. A0150712052 (J.-P.P.) and grant GC010815453 (Grand Challenge H100 Jean Zay, J.-P.P.).